OpenStudy (dmezzullo):

Which of the following equations is written in slope-intercept form? y - 4x = -3 y = 4x - 3 -4x = -y - 3 x = one fifth(y + 3)

OpenStudy (anonymous):

slope intercept form is y=mx+c, so it is the second one.

OpenStudy (anonymous):

explain it to him like the steps lol plz thats what he wants

OpenStudy (anonymous):

*y=mx+b* and its the second one.

OpenStudy (anonymous):

-4=m -3=b

OpenStudy (anonymous):

There are no steps, questions of form are a definition, no steps involved

OpenStudy (anonymous):

lol i read it wrong sorry nvm

OpenStudy (dmezzullo):

tyvm

OpenStudy (anonymous):

tf does that mean? ^ lol

OpenStudy (anonymous):

thank u vary much

OpenStudy (anonymous):

SMH ohh lol